How long it will take a wave travelling from one end of 85 m long and 5 kg wire to the other end if the tension of 300 N is applied?
1
Expert's answer
2019-10-22T10:03:55-0400

The speed of the wave


"v=\sqrt{\frac{T}{\mu}}=\sqrt{\frac{T}{m/l}}=\sqrt{\frac{300N}{5kg/85m}}=71.4\: m/s"

The time of a wave traveling


"t=\frac{l}{v}=\frac{85\: m}{71.4\: m/s}=1.2\: s"